OpenStudy (anonymous):

Factor. 4x2 + 10x A. 2x(2x + 5) B. x(2x + 5) C. 4x(x + 5) D. 2x2(2x + 5)

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

\[\huge 4x^2 + 10x = 2x( 2x+5)\] Hence option A is correct one

OpenStudy (anonymous):

\(4x^2 + 10x = 2x(2x + 5)\) because \(ab + ab = a(b + c)\)

OpenStudy (anonymous):

In this case, a = 2x, b = 2x and c = 5.

OpenStudy (anonymous):

so is it C

OpenStudy (anonymous):

oh nvm its A